OpenStudy (anonymous):

Three boxes in contact rest side-by-side on a smooth, horizontal floor. Their masses are 5.0 kg, 3.0 kg, and 2.0 kg, with the 3.0 kg box in the center. A force of 50 N pushes on the 5.0 kg box, which pushes on the other two boxes. What magnitude of force does the 3.0 kg box exert on the 5.0 kg box? What magnitude of force does the 3.0 kg box exert on the 2.0 kg box?

OpenStudy (michele_laino):

please, apply the third principle of Newton, or principle of "action-reaction"

OpenStudy (fretje):

not enough information, it seems. Does the floor with its smooth surface exert a shear force on the boxes? Do the boxes keep resting, or do they accelerate?

OpenStudy (fretje):

i found it: Three boxes rest side-by-side on a smooth, horizontal floor. Their masses are 5.0 kg, 3.0 kg, and 2.0 kg, with the 3.0-kg mass in the center. A force of 50 N pushes on the 5.0-kg mass, which pushes against the other two masses. What is the contact force between the 3.0-kg mass and the 2.0-kg mass? Choices: 0 N 10 N 25 N 40 N

OpenStudy (fretje):

since the masses are side by side, and move in unison Fnet = ma 50 N = (5+3+2)a 50 N = (10 kg) a a = 50/10 a = 5 m/s^2 The Contact force between the 3 kg and 2 kg The net Force on 2 kg block, (Fnet2) is the Force from 3 kg block (F32) and is equal to ma Fnet2 = F32 = ma F32 = (2 kg)(5 m/s^2) = 10 N F23 = F32 is the reaction of the 2 kg block on the 3 kg block the answer is 10 N

OpenStudy (fretje):

i copied this from yahoo https://answers.yahoo.com/question/index?qid=20081105130816AAQKokS

OpenStudy (abhisar):